Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Veritas Probability Explanation required

Expert replies
Source: — Problem Solving |

by LalaB » Sun May 20, 2012 9:12 pm
to be accepted at least one school =1- not to be accepted at all


not to be accepted at all =(4/5)*(4/5)*(4/5)=64/125

1-64/125=61/125
Happy are those who dream dreams and are ready to pay the price to make them come true.(c)

In order to succeed, your desire for success should be greater than your fear of failure.(c)
Join the discussion

by karthikpandian19 » Sun May 20, 2012 9:33 pm
To be accepted by atleast one school is also equal to OR condition, so why cant it be 1/5 +1/5 + 1/5 = 3/5 ???

Please clarify the confusion???
LalaB wrote:to be accepted at least one school =1- not to be accepted at all


not to be accepted at all =(4/5)*(4/5)*(4/5)=64/125

1-64/125=61/125
Join the discussion

by LalaB » Sun May 20, 2012 9:49 pm
To be accepted by atleast one school means (1 -to be accepted to NONE school at all)

if ur get OR condition, then u wont get NONE, since Or means that there is some possibibity for being accepted
Happy are those who dream dreams and are ready to pay the price to make them come true.(c)

In order to succeed, your desire for success should be greater than your fear of failure.(c)
Join the discussion

by diebeatsthegmat » Sun May 20, 2012 10:28 pm
karthikpandian19 wrote:Image
this one is quite easy
step 1: one of 3 schools he will get in :
3!/2!* 20/100*80/100*80/100=48/125
step 2: he will get in 2 in three:
3!/2!*20/100*20/100*80/100=12/125
step 3: he is able to get in all the whole three
3!/3!*20/100*20/100*20/100=1/125
totally =(48+12+1)/125= the answwer
Join the discussion

by karthikpandian19 » Sun May 20, 2012 10:46 pm
Can any GMAT Expert explain this one?

Still i m confused????
karthikpandian19 wrote:Image
Join the discussion

by aneesh.kg » Fri Jun 01, 2012 9:46 pm
karthikpandian19 wrote:Can any GMAT Expert explain this one?

Still i m confused????
Ok. Let me try.
Remember: AND implies 'multiplication'. OR implies 'addition'.

Let the names of the three MBA programs be A, B and C.
If for each of A, B, C
P(accepted) = 1/5, then
for each of them
P(not accepted) = 1 - 1/5 = 4/5

Better Method:
P(accepted by atleast one school)
= 1 - P(not accepted by any school)
= 1 - P(not accepted by A) AND P(not accepted by B) AND P(not accepted by C)
= 1 - (4/5)*(4/5)*(4/5)
= 1 - (64/125)
= [spoiler]61/125[/spoiler]

Alternatively(And, this is a long method):

I think you were stuck in the explanation of this method, so I will elaborate it in detail. Read on.

P(accepted by atleast one school)
= P(accepted by one school) OR P(accepted by two schools) OR P(accepted by three schools)

P(accepted by one school)
= [P(accepted by A) AND P(not accepted by B) AND P(not accepted by C)]*[Arrange]
Why and What do we arrange?
Notice that we have considered a Yes-No-No for A,B,C but we can also have a No-Yes-No or a No-No-Yes.
The Yes,No,No can be arranged in (3!/2!) = 3 ways.
So,
P(accepted by one school)
= (1/5)*(4/5)*(4/5)*3
= 48/125

P(accepted by two schools)
= [P(accepted by A) AND P(accepted by B) AND P(not accepted by C)]*[Arrange]
= (1/5)*(1/5)*(4/5)*3
= 12/125

Note that we have arranged a Yes-Yes-No in 3!/2! = 3 ways here.
P(accepted by 3 schools)
= [P(accepted by A) AND P(accepted by B) AND P(accepted by C)]*[Arrange]
= (1/5)*(1/5)*(1/5)*1
= 1/125

Note that there is just one way of arranging Yes-Yes-Yes.
Therefore,
P(accepted by atleast one school)
= (48/125) + (12/125) + (1/125)
= [spoiler]61/125[/spoiler]

Same Answer. Hooraah!
Aneesh Bangia
GMAT Math Coach
[email protected]

GMATPad:
Facebook Page: https://www.facebook.com/GMATPad
Join the discussion